OpenStudy (anonymous):

Find the perimeter of the rectangle pictured. A rectangle with a length of 9x minus 5 and a width of 2 x + 4 A. 11x + 9 B. 22x - 2 C. 18x - 20 D. 11x - 1

OpenStudy (anonymous):

\[\left( 9x+4 \right)\left( 2x-5 \right)\]\[=18x ^{2}+26x-20\]

OpenStudy (anonymous):

oops read it wrong…sorry

OpenStudy (anonymous):

the answer is still right I just switched the +4 and -5 when I typed it

OpenStudy (anonymous):

nevermind…i did read it wrong I'm trying to find area not perimeter haha

OpenStudy (anonymous):

2x+4+2x+4+9x-5+9x-5 =22x-2 B